OpenStudy (anonymous):

conjugate of a real no is again that specific real no then why conjugate of 2+sqrt3 is 2-sqrt 3???

Parth (parthkohli):

I don't get your question since \(2 - \sqrt{3}\) is a real number as well.

OpenStudy (anonymous):

no i mean conjugate of a real no is again that real no means x conjugate is x if x

OpenStudy (anonymous):

belongs to R

Parth (parthkohli):

\(2 + \sqrt{3}\) and \(2 - \sqrt{3}\) both belong to \(\mathbb R\).

OpenStudy (anonymous):

but both are not same

Parth (parthkohli):

Eh? Two conjugates are never the same... In fact, two unequal real numbers are never the same

OpenStudy (anonymous):

in complex analysis it is a result that "conjugate of z is z iff z is real" get it?

Parth (parthkohli):

But this thing is talking about the complex conjugate. :-)

Parth (parthkohli):

The general conjugate and the complex conjugate are totally different.

Parth (parthkohli):

The complex conjugate of \(2 + \sqrt{3}\) is \(2 + \sqrt{3}\) only. But the general conjugate has no explicit definition. It's just changing the sign in between.

OpenStudy (anonymous):

whats the difference.. then what do you say about conjugate of 2???

OpenStudy (anonymous):

i dont think there may be a difference between real and complex conjugate